Eliminate the parameter t from the following. z(t) = 2 sint - 6 ly(t) = 3 cost - 6 Preview
Eliminate the parameter t from the following. S *(t) = sint - 5 ly(t) = cost + 3 Preview
Eliminate the parameter t from the following. { r(t) = 4 + 4 sint y(t) = 2 + 2 cost Preview
Eliminate the parameter t from the following. *(t) = 5 tant 1 y(t) = 5 sect Preview
Eliminate the parameter t from the following. S (t) = 6 + 2 tant 1 y(t) = 2 + 2 sect Preview
Answer #1

Eliminate the parameter t from the following set 2 sint - 6 Lyle 3 cost-6 A the given 2(t Үt) a sint - 6 3 cost – s 2 2 sintEliminate the parameter t from the following alt sint - 5 ༽ ༧༤ cost +3 the given parametert from 지 sint - 5 cost + 3 3 [ Additherefore a(t) = 4+4 sint 3) =) x-4 = 4 sin(t) (x-4) A 4 sint sin²t yet [ Subtoactingoy 4 both side [divided by 4 both side]a = stant tant u by s both side =) tant [ dividing by * ? [ Squaring both side] (I) And st a 5 seet 2) y sect [ dividing byx-6 2 tant =) tant =) X-6 A (I ) 7 R-C [ dididing both side by z] tant Squaring both side tant () - And afft y-2= aseet I sub